A. Determine the magnitude of the force on an electron traveling 5.40×105 m/s horizontally to the east in a vertically upward magnetic field of strength 0.33 T .

B. Determine the direction of the force on an electron.
Answer #1

a)

here

Fmax = q * v * B

Fmax = 1.6 * 10^-19 * 5.4 * 10^5 * 0.33

Fmax = 2.85 * 10^-14 N

b)

by the right hand rule the force must be directed to the north